OpenStudy (anonymous):

Solve: 5(x − 1) − (x + 5) = 2

OpenStudy (anonymous):

5x - 5 - x - 5 = 2 4x=12. i think rest u can solve

OpenStudy (anonymous):

thanks. what about this problem

OpenStudy (anonymous):

x-5 over 4 = 3 solve for x.

OpenStudy (anonymous):

x-5=12.

OpenStudy (anonymous):

just shift 4 the other side n solve for x

OpenStudy (anonymous):

3 + x = 4 is the same as 4 = 3 + x. <is that associative property?

OpenStudy (anonymous):

yes right..

OpenStudy (anonymous):

okay making sure. thank you